# Investigating photometric and spectroscopic variability in the multiply-imaged ‘Little Red Dot’ A2744-QSO1

### v2

@@
JWST observations have uncovered a new population of red, compact objects at high redshifts dubbed ‘Little Red Dots’ (LRDs), which typically show broad emission lines and are thought to be dusty Active Galactic Nuclei (AGN). Some of their other features, however, challenge the AGN explanation, such as prominent Balmer breaks and extremely faint or even missing metal high-ionization lines, X-ray, or radio emission, including in deep stacks. Time variability is a unique test of AGN activity. Here, we exploit the multiply-imaged LRD A2744-QSO1, which offers a unique test of variability due to lensing-induced time delays between the three images spanning 22 yr (2.7 yr in the rest-frame), to investigate its photometric and spectroscopic variability. We find the equivalent widths (EWs) of the broad H and H lines, which are independent of magnification and other systematics, to exhibit significant variations, up to % for H and up to % in H, on a timescale of 875 d (2.4 yr) in the rest-frame. This suggests that A2744-QSO1 is indeed an AGN. We find no significant photometric variability beyond the limiting systematic uncertainties, so it currently cannot be determined whether the EW variations are due to line-flux or continuum variability. These results are consistent with a typical damped random walk (DRW) variability model for an AGN like A2744-QSO1 () given the sparse sampling of the light-curve with the available data. Our results therefore support the AGN interpretation of this LRD, and highlight the need for further photometric and spectroscopic monitoring in order to build a detailed and reliable light-curve.
@@
One of the most prominent results from the JWST (Gardner et al., 2023; McElwain et al., 2023) in its first few years of operations is the serendipitous discovery of a new population of extremely red point-sources at high redshifts (), the so-called ‘Little Red Dots’ (LRDs; e.g. Matthee et al., 2024a). These were initially identified in JWST imaging data (e.g. Endsley et al., 2023; Furtak et al., 2023a; Barro et al., 2024; Labbé et al., 2025) with follow-up later revealing broad emission lines, consistent with heavily dust-attenuated (), broad-line type 1 active galactic nuclei (AGN; e.g. Kocevski et al., 2023; Harikane et al., 2023; Matthee et al., 2024a; Furtak et al., 2024; Kokorev et al., 2023; Greene et al., 2024; Killi et al., 2023), although the picture has grown more complicated with time (e.g. Pérez-González et al., 2024; Baggen et al., 2024). Many LRD AGN have been observed to date, down to (e.g. Wang et al., 2024a; Kocevski et al., 2024; Naidu et al., 2024), with the largest photometric samples encompassing several hundred objects selected in JWST’s blank fields (e.g. Kokorev et al., 2024; Kocevski et al., 2024). At , LRDs seem to represent 20 % to 30 % of the AGN population (e.g. Kocevski et al., 2024). If LRDs are indeed powered by an accreting black hole, they pose a challenge for black hole growth models as they present extreme black-hole-to-galaxy mass ratios (e.g. Kokorev et al., 2023; Furtak et al., 2024; Durodola et al., 2024; Matthee et al., 2024b) which are not observed in the low-redshift Universe (though cf. Lin et al. 2024).
@@
Most LRDs are, however, extremely faint in X-rays (e.g. Ananna et al., 2024; Yue et al., 2024) or radio (Perger et al., 2024; Mazzolari et al., 2024; Gloudemans et al., 2025), and seem to lack a hot dust component in the mid-infrared (e.g. Williams et al., 2024; Akins et al., 2024). These features could be explained with e.g. Compton-thick broad-line regions (Maiolino et al., 2024), super-Eddington accretion (e.g. Pacucci & Narayan, 2024; Lambrides et al., 2024; Inayoshi et al., 2024; Madau & Haardt, 2024; Madau, 2025), or seeds of primordial origin (e.g. Dayal, 2024). However, they have also, in combination with the detection of distinct Balmer-breaks in some LRDs (e.g. Wang et al., 2024b; Setton et al., 2024; Ma et al., 2024; Labbé et al., 2024), led to alternative models where their spectral energy distribution (SED) is dominated by stars (e.g. Pérez-González et al., 2024; Wang et al., 2024b). While the extreme stellar densities implied in such a picture could a priori help explain the observed broad lines (Baggen et al., 2024), such extreme densities are not otherwise observed in the Universe. On the other hand, the presence of a Balmer-break does not necessarily contradict an AGN interpretation, as material around the black hole can in theory have physical conditions such that a significant fraction of the hydrogen atoms have their electrons in or above the second energy level (e.g. Laor & Davis, 2011; Inayoshi & Maiolino, 2024; Li et al., 2024; Ji et al., 2025).
@@
Since all type 1 AGN vary temporally (e.g. Matthews & Sandage, 1963; Kelly et al., 2009; MacLeod et al., 2012; Cammelli et al., 2025), variability could be a key discriminator between AGN and stellar-only models. Recent studies have therefore also looked for evidence of variability in LRDs (Kokubo & Harikane, 2024; Zhang et al., 2024; Tee et al., 2024; Ji et al., 2025). The investigated samples show mostly no variability in several epochs of broad-band photometry, both in the rest-frame ultra-violet (UV) and optical (Kokubo & Harikane, 2024; Zhang et al., 2024; Tee et al., 2024) with only a small subsample of LRDs displaying potential variability (Zhang et al., 2024). However, since JWST has only been operational for a little over two years, time-domain data are still sparse. Therefore, variability studies of LRDs so far are based on a handful of JWST epochs or rely on archival Hubble Space Telescope (HST) data (e.g. Tee et al., 2024), which only probe the faint rest-frame UV emission of LRDs.
@@
In this work, we examine the variability of the multiply-imaged LRD A2744-QSO1 (Furtak et al., 2023a, 2024), one of the first LRDs ever detected. It was first identified photometrically as a uniquely red-and-compact object in the Ultra-deep NIRSpec and NIRCam ObserVations before the Epoch of Reionization (UNCOVER; Bezanson et al., 2022) observations of the strong lensing (SL) cluster Abell 2744 (A2744; e.g. Abell et al., 1989; Merten et al., 2011) and then spectroscopically confirmed as a broad-line LRD at (Furtak et al., 2023a, 2024). Being multiply-imaged, A2744-QSO1 is ideal for this type of study because SL induces a delay in the relative arrival time of each of the three images. This means that each visit effectively provides three epochs of observation, separated by several years. With two years of JWST imaging and spectroscopy of A2744-QSO1 available to date, we can investigate it for variability in its emission over a two-decade timescale (see Ji et al. (2025) for another recent analysis of variability in this object; we also compare to their results in our discussion).
@@ 2 Data
In the two years since the first JWST observations of UNCOVER, the center of A2744 was imaged numerous times with JWST. Here, we use the Grism redshift & line analysis software for space-based slitless spectroscopy (grizli; Brammer, 2023a) to query the Mikulski Archive for Space Telescopes (MAST) for Near Infrared Camera (NIRCam; Rieke et al., 2023) imaging targeted at the cluster center, and to perform the data reduction. Each filter and epoch is drizzled into mosaics of 0.04″ per pixel. We list our six imaging epochs in Tab. 2. The NIRCam observations reduced here include data from programs 2756 (PI: W. Chen), 2561 (UNCOVER; PIs: I. Labbé & R. Bezanson; Bezanson et al., 2022), 2883 (Magnif; PI: F. Sun), 3538 (PI: E. Iani), 4111 (MegaScience; PI: K. Suess; Suess et al., 2024), and 3516 (All the Little Things; PIs: R. Naidu & J. Matthee; Naidu et al., 2024, ALT).
Our spectroscopic analysis is based on data from the UNCOVER program (Price et al., 2024), which also obtained ultra-deep Near Infrared Spectrograph (NIRSpec; Jakobsen et al., 2022; Böker et al., 2023) prism observations using the micro-shutter array (MSA; Ferruit et al., 2022; Jakobsen, 2024) in July 2023. Due to a guide-star acquisition failure in the initial observations, a second epoch of prism spectra was taken in July 2024, which includes image A of A2744-QSO1 in the MSA. For this work, we utilize v4 reductions from the DAWN JWST Archive (DJA; Heintz et al., 2024; de Graaff et al., 2024), which are reduced using the development version (v0.9.5) of MSAEXP (Brammer, 2023b). We refer the reader to Price et al. (2024) for most of the details of the MSAEXP reduction and refer to appendix B for more information. The primary improvement of the new reductions is a substantial extension of the wavelength coverage to include the full NIRSpec prism disperser throughput to cover wavelengths , achieving spectral resolutions . This has the advantage of fully covering the H line for our target on the red end of the detector (see section 3). While several JWST grism spectroscopy programs targeted A2744 as well, namely GLASS-JWST (ID 1324; Treu et al., 2022), Magnif, and ALT, GLASS-JWST does not cover wavelength ranges with emission lines for our target, and Magnif only covers the H line which has too low a signal-to-noise ratio (SNR) to be exploited here. The ALT spectroscopy covers the H line, but only shows marginal detections in two images and is therefore not usable for our variability studies.
@@ 3 Spectroscopic analysis
We measure the line fluxes by fitting spectra as explained in detail in appendix B. The resulting integrated emission line fluxes, underlying continua, and rest-frame equivalent-widths (EWs) are listed in Tab. 3 for each image and epoch. Unlike absolute line fluxes, the EWs are neither prone to lensing nor instrumental systematics (e.g. slit-losses), which is why we use them as the primary and most robust indicators of variability (see also Ji et al., 2025). As can be seen in Fig. 2, both H and H show a consistent variability in their EW over time222Here and throughout this letter, we define the time axis in days relative to the UNCOVER imaging (epoch 2 at , see Tab. 2). The relative time is then obtained by adding the difference between the epochs and , the gravitational time delay listed in Tab. 1.. More quantitatively, we note a drop in EW of % in H and % in H between the times probed by image C and image A (in both its epochs). The absolute calibration of the red edge of the NIRSpec prism, where H is located, is uncertain ( %), but this should not impact the EW measurements directly.
@@ 5 Discussion and conclusion
Our variability measurements in the multiply-imaged LRD A2744-QSO1 come after several years of debate as to the nature of LRDs, which started when they were first identified photometrically (e.g. Furtak et al., 2023a; Barro et al., 2024; Labbé et al., 2025), and has continued as more and more of the puzzling aspects of their SEDs were revealed (e.g. Pérez-González et al., 2024; Ananna et al., 2024; Yue et al., 2024; Setton et al., 2024; Gloudemans et al., 2025). Because of the elusive nature of the LRDs’ SED (e.g. Setton et al., 2024; Labbé et al., 2024), at this stage we cannot predict which part of the SED – the optical or the UV continuum – can be expected to vary in the first place. The hardest component of the SED to explain with a purely stellar model are the broad emission lines – both their high luminosity and their width. And indeed, we do find these to vary consistently in H and H, by % and %, respectively as shown in section 3. Note that Ji et al. (2025) find a similar result for H in the same source, which we here confirm via the higher SNR H line. The current data, however, are insufficient to securely determine whether the EW variations are dominated by variations in the continuum, the line fluxes, or perhaps most probably – both. If dominated by continuum variation, our detection of broad-line EW variability in section 3 would correspond to magnitude excursions, which are below the limiting sensitivity of our photometric analysis in section 4.1.
@@
While the last year in particular saw several studies looking into photometric variability of LRDs, the vast majority of LRDs did not display significant variations over the handful of epochs considered (Kokubo & Harikane, 2024; Zhang et al., 2024; Tee et al., 2024), similar to what we find for each image individually. Only eight out of several hundred investigated LRDs seem to present significant variability (Zhang et al., 2024). This could a priori indicate the LRD population at large to not necessarily be AGN, although it has also been claimed that e.g. super-Eddington accretion could explain the lack of variability (e.g. Inayoshi et al., 2024). In the context of these non-detections in the literature, our photometric results do not come as a surprise. Even when combining the three images, and thus tripling the number of effective epochs to , we do not find any significant photometric variability, i.e. large magnitude excursions (see section 4.1). This is nonetheless still consistent with the stochastic nature of AGN variability (e.g. Kelly et al., 2009; MacLeod et al., 2010), as we have shown in section 4.2. Indeed, our DRW simulations show that even if there were intrinsic photometric variability in A2744-QSO1, the probability of detecting it with our sampling is low and nearly impossible in the small amount of epochs for each image separately. The lack of significant photometric variability in LRDs in the literature therefore cannot rule out intrinsic variability of LRDs, nor that they are AGN. It merely means that with the data available so far, the light-curves of LRDs are not sufficiently sampled to rule out expected levels of photometric variability.
@@
In conclusion, our detection of significant broad-line EW variability in the multiply-imaged LRD A2744-QSO1 presents strong evidence in support of the AGN interpretation of this object, and a varying AGN at that, even if we do not find significant photometric variability at this stage. Our results show that robust investigations into the variability of LRDs, including A2744-QSO1, will require dedicated monitoring with numbers of epochs large enough to sample the light-curve frequently and maximize the probability to detect significant variation events. More importantly, we stress that such monitoring campaigns need to be consistent in depth, coverage, and observing strategy to minimize the systematics between single epochs. Thanks to the gravitational time delays, this source is ideally suited for further study of its variability and could make for a prime target for a large-scale reverberation mapping campaign in the future. With that, one could potentially measure the black-hole mass of an AGN at directly and, for the first time, calibrate the scaling relations used to measure black-hole masses beyond the near-by Universe (e.g. Liu et al., 2024). Thanks to SL time delays, multiply-imaged, cluster-lensed AGN like A2744-QSO1 may open a door to achieve this on a reasonable timescale out to high-redshifts and high black-hole masses (e.g. Golubchik et al. 2024).

### v3

@@
JWST observations have uncovered a new population of red, compact objects at high redshifts dubbed ‘Little Red Dots’ (LRDs), which typically show broad emission lines and are thought to be dusty Active Galactic Nuclei (AGN). Some of their other features, however, challenge the AGN explanation, such as prominent Balmer breaks and extremely faint or even missing metal high-ionization lines, X-ray, or radio emission, including in deep stacks. Time variability is another, robust, test of AGN activity. Here, we exploit the multiply-imaged LRD A2744-QSO1, which offers a particularly unique test of variability due to lensing-induced time delays between the three images spanning 22 yr (2.7 yr in the rest-frame), to investigate its photometric and spectroscopic variability. We find the equivalent widths (EWs) of the broad H and H lines, which are independent of magnification and other systematics, to exhibit significant variations, up to % for H and up to % in H, on a timescale of 875 d (2.4 yr) in the rest-frame. This suggests that A2744-QSO1 is indeed an AGN. We find no significant photometric variability beyond the limiting systematic uncertainties, so it currently cannot be determined whether the EW variations are due to line-flux or continuum variability. These results are consistent with a typical damped random walk (DRW) variability model for an AGN like A2744-QSO1 () given the sparse sampling of the light-curve with the available data. Our results therefore support the AGN interpretation of this LRD, and highlight the need for further photometric and spectroscopic monitoring in order to build a detailed and reliable light-curve.
@@
One of the most prominent results from the JWST (Gardner et al., 2023; McElwain et al., 2023) in its first few years of operations is the serendipitous discovery of a new population of extremely red point-sources at high redshifts (), the so-called ‘Little Red Dots’ (LRDs; e.g. Matthee et al., 2024a). These were initially identified in JWST imaging data (e.g. Endsley et al., 2023; Labbé et al., 2023, 2025; Furtak et al., 2023a; Barro et al., 2024) with follow-up spectroscopy later revealing broad emission lines, consistent with heavily dust-attenuated (), broad-line type 1 active galactic nuclei (AGN; e.g. Kocevski et al., 2023; Harikane et al., 2023; Matthee et al., 2024a; Furtak et al., 2024; Kokorev et al., 2023; Greene et al., 2024; Killi et al., 2024), although the picture has grown more complicated with time (e.g. Pérez-González et al., 2024; Baggen et al., 2024, 2025). Many LRD AGN have been observed to date, down to (e.g. Wang et al., 2024a; Kocevski et al., 2024; Naidu et al., 2024), with the largest photometric samples encompassing several hundred objects selected in JWST’s blank fields (e.g. Kokorev et al., 2024; Kocevski et al., 2024). At , LRDs seem to represent 20 % to 30 % of the AGN population (e.g. Kocevski et al., 2024) and tend to reside in over-densities (e.g. Schindler et al., 2024; Matthee et al., 2024b). If LRDs are indeed powered by an accreting black hole, they pose a challenge for black hole growth models as they present extreme black-hole-to-galaxy mass ratios (e.g. Kokorev et al., 2023; Furtak et al., 2024; Durodola et al., 2024; Matthee et al., 2024b) which are not observed in the low-redshift Universe (though cf. Lin et al. 2025; Euclid Collaboration et al. 2025).
@@
Most LRDs are, however, extremely faint in X-rays (e.g. Ananna et al., 2024; Yue et al., 2024) or radio (Mazzolari et al., 2024; Perger et al., 2025; Gloudemans et al., 2025), and seem to lack a hot dust component in the mid-infrared (e.g. Williams et al., 2024; Akins et al., 2024; Setton et al., 2025). These features could be explained with e.g. Compton-thick broad-line regions (Maiolino et al., 2025), super-Eddington accretion (e.g. Pacucci & Narayan, 2024; Lambrides et al., 2024; Inayoshi et al., 2024; Madau & Haardt, 2024; Madau, 2025; Tripodi et al., 2024), or seeds of primordial origin (e.g. Dayal, 2024). However, they have also, in combination with the detection of distinct Balmer-breaks in some LRDs (e.g. Wang et al., 2024b; Setton et al., 2024; Ma et al., 2025; Labbé et al., 2024), led to alternative models where their spectral energy distribution (SED) is dominated by stars (e.g. Pérez-González et al., 2024; Wang et al., 2024b). While the extreme stellar densities implied in such a picture could a priori help explain the observed broad lines (Baggen et al., 2024, 2025), such extreme densities are not otherwise observed in the Universe. On the other hand, the presence of a Balmer-break does not necessarily contradict an AGN interpretation, as material around the black hole can in theory have physical conditions such that a significant fraction of the hydrogen atoms have their electrons in or above the second energy level (e.g. Laor & Davis, 2011; Inayoshi & Maiolino, 2025; Li et al., 2025; Ji et al., 2025; Naidu et al., 2025).
@@
Since all type 1 AGN vary temporally (e.g. Matthews & Sandage, 1963; Kelly et al., 2009; MacLeod et al., 2012; Cammelli et al., 2025), variability could be a key discriminator between AGN and stellar-only models. Recent studies have therefore also looked for evidence of variability in LRDs (Kokubo & Harikane, 2024; Zhang et al., 2024; Tee et al., 2025; Ji et al., 2025). The investigated samples show mostly no variability in several epochs of broad-band photometry, both in the rest-frame ultra-violet (UV) and optical (Kokubo & Harikane, 2024; Zhang et al., 2024; Tee et al., 2025) with only a small subsample of LRDs displaying potential variability (Zhang et al., 2024). However, since JWST has only been operational for a little over two years, time-domain data are still sparse. Therefore, variability studies of LRDs so far are based on a handful of JWST epochs or rely on archival Hubble Space Telescope (HST) data (e.g. Tee et al., 2025), which only probe the faint rest-frame UV emission of LRDs.
@@
In this work, we examine the variability of the multiply-imaged LRD A2744-QSO1 (Furtak et al., 2023a, 2024), one of the first LRDs ever detected. It was first identified photometrically as a uniquely red-and-compact object in the Ultra-deep NIRSpec and NIRCam ObserVations before the Epoch of Reionization (UNCOVER; Bezanson et al., 2024) observations of the strong lensing (SL) cluster Abell 2744 (A2744; e.g. Abell et al., 1989; Merten et al., 2011) and then spectroscopically confirmed as a broad-line LRD at (Furtak et al., 2023a, 2024). Being multiply-imaged, A2744-QSO1 is ideal for this type of study because SL induces a delay in the relative arrival time of each of the three images. This means that each visit effectively provides three epochs of observation, separated by several years. With two years of JWST imaging and spectroscopy of A2744-QSO1 available to date, we can investigate it for variability in its emission over a two-decade timescale (see Ji et al. (2025) for another recent analysis of variability in this object; we also compare to their results in our discussion).
@@ 2 Data
In the two years since the first JWST observations of UNCOVER, the center of A2744 was imaged numerous times with JWST. Here, we use the Grism redshift & line analysis software for space-based slitless spectroscopy (grizli; Brammer, 2023a) to query the Mikulski Archive for Space Telescopes (MAST) for Near Infrared Camera (NIRCam; Rieke et al., 2023) imaging targeted at the cluster center, and to perform the data reduction. Each filter and epoch is drizzled into mosaics of 0.04″ per pixel. We list our six imaging epochs in Tab. 2. The NIRCam observations reduced here include data from programs 2756 (PI: W. Chen), 2561 (UNCOVER; PIs: I. Labbé & R. Bezanson; Bezanson et al., 2024), 2883 (Magnif; PI: F. Sun), 3538 (PI: E. Iani), 4111 (MegaScience; PI: K. Suess; Suess et al., 2024), and 3516 (All the Little Things; PIs: R. Naidu & J. Matthee; Naidu et al., 2024, ALT).
Our spectroscopic analysis is based on data from the UNCOVER program (Price et al., 2025), which also obtained ultra-deep Near Infrared Spectrograph (NIRSpec; Jakobsen et al., 2022; Böker et al., 2023) prism observations using the micro-shutter array (MSA; Ferruit et al., 2022; Jakobsen, 2024) in July 2023. Due to a guide-star acquisition failure in the initial observations, a second epoch of prism spectra was taken in July 2024, which includes image A of A2744-QSO1 in the MSA. For this work, we utilize v4 reductions from the DAWN JWST Archive (DJA; Heintz et al., 2024; de Graaff et al., 2024a), which are reduced using the development version (v0.9.5) of MSAEXP (Brammer, 2023b). We refer the reader to Price et al. (2025) for most of the details of the MSAEXP reduction and refer to appendix B for more information. The primary improvement of the new reductions is a substantial extension of the wavelength coverage to include the full NIRSpec prism disperser throughput to cover wavelengths , achieving spectral resolutions . This has the advantage of fully covering the H line for our target on the red end of the detector (see section 3). While several JWST grism spectroscopy programs targeted A2744 as well, namely GLASS-JWST (ID 1324; Treu et al., 2022), Magnif, and ALT, GLASS-JWST does not cover wavelength ranges with emission lines for our target, and Magnif only covers the H line which has too low a signal-to-noise ratio (SNR) to be exploited here. The ALT spectroscopy covers the H line, but only shows marginal detections in two images and is therefore not usable for our variability studies.
@@ 3 Spectroscopic analysis
– Note: The line fluxes and underlying continuum flux densities are de-magnified, but the EWs are independent of magnification.
@@
We measure the line fluxes by fitting spectra as explained in detail in appendix B. The resulting integrated emission line fluxes, underlying continua, and rest-frame equivalent-widths (EWs) are listed in Tab. 3 for each image and epoch. Unlike absolute line fluxes, the EWs are neither prone to lensing nor instrumental systematics (e.g. slit-losses), which is why we use them as the primary and most robust indicators of variability (see also Ji et al., 2025). As can be seen in Fig. 2, both H and H show a consistent variability in their EW over time222Here and throughout this letter, we define the time axis in days relative to the UNCOVER imaging (epoch 2 at , see Tab. 2). The relative time is then obtained by adding the difference between the epochs and , the gravitational time delay listed in Tab. 1.. More quantitatively, we note a drop in EW of % in H and % in H between the times probed by image C and image A (in both its epochs). The absolute calibration of the red edge of the NIRSpec prism, where H is located, is uncertain ( %), but this should not impact the EW measurements directly. Finally, we do not find any variability in the broad line widths within the uncertainties.
@@ 5 Discussion and conclusion
Our variability measurements in the multiply-imaged LRD A2744-QSO1 come after several years of debate as to the nature of LRDs, which started when they were first identified photometrically (e.g. Labbé et al., 2023, 2025; Furtak et al., 2023a; Barro et al., 2024), and has continued as more and more of the puzzling aspects of their SEDs were revealed (e.g. Pérez-González et al., 2024; Ananna et al., 2024; Yue et al., 2024; Setton et al., 2024; Gloudemans et al., 2025). Because of the elusive nature of the LRDs’ SED (e.g. Setton et al., 2024; Labbé et al., 2024), at this stage we cannot predict which part of the SED – the optical or the UV continuum – can be expected to vary in the first place. The hardest component of the SED to explain with a purely stellar model are the broad emission lines – both their high luminosity and their width. And indeed, we do find these to vary consistently in H and H, by % and %, respectively as shown in section 3. Note that Ji et al. (2025) find a similar result for H in the same source, which we here confirm via the higher SNR H line. The current data, however, are insufficient to securely determine whether the EW variations are dominated by variations in the continuum, the line fluxes, or perhaps most probably – both. If dominated by continuum variation, our detection of broad-line EW variability in section 3 would correspond to magnitude excursions, which are below the limiting sensitivity of our photometric analysis in section 4.1.
@@
While the last year in particular saw several studies looking into photometric variability of LRDs, the vast majority of LRDs did not display significant variations over the handful of epochs considered (Kokubo & Harikane, 2024; Zhang et al., 2024; Tee et al., 2025), similar to what we find for each image individually. Only eight out of several hundred investigated LRDs seem to present significant variability (Zhang et al., 2024). This could a priori indicate the LRD population at large to not necessarily be AGN, although it has also been claimed that e.g. super-Eddington accretion could explain the lack of variability (e.g. Inayoshi et al., 2024). In the context of these non-detections in the literature, our photometric results do not come as a surprise. Even when combining the three images, and thus tripling the number of effective epochs to , we do not find any significant photometric variability, i.e. large magnitude excursions (see section 4.1). This is nonetheless still consistent with the stochastic nature of AGN variability (e.g. Kelly et al., 2009; MacLeod et al., 2010), as we have shown in section 4.2. Indeed, our DRW simulations show that even if there were intrinsic photometric variability in A2744-QSO1, the probability of detecting it with our sampling is low and nearly impossible in the small amount of epochs for each image separately. The lack of significant photometric variability in LRDs in the literature therefore cannot rule out intrinsic variability of LRDs, nor that they are AGN. It merely means that with the data available so far, the light-curves of LRDs are not sufficiently sampled to rule out expected levels of photometric variability.
Our detection of significant broad-line EW variability in the multiply-imaged LRD A2744-QSO1 presents strong evidence in support of the AGN interpretation of this object, and a varying AGN at that, even if we do not find significant photometric variability at this stage. At the time of writing, high-resolution JWST/NIRSpec spectroscopy of A2744-QSO1 has revealed strong narrow absorption features in the Balmer-lines (Ji et al., 2025; D’Eugenio et al., 2025), which further corroborate the scenario where this object is dominated by an AGN, though of lower black-hole mass than previous estimates, with a nearly absent host galaxy (; D’Eugenio et al., 2025). Most recently, the detection of two peculiar LRD objects, with extreme Balmer-breaks and a complete absence of metal lines as in A2477-QSO1 (Naidu et al., 2025; de Graaff et al., 2025), has further shed light on the nature of LRDs: The extreme Balmer-break, broad lines, and absorption features in said lines are consistent with emission from a super-massive black hole embedded in an extremely dense cocoon of hot gas, similar to a stellar atmosphere (e.g. Inayoshi & Maiolino, 2025; Ji et al., 2025; Naidu et al., 2025), which in addition suppresses the radio and X-ray emission (e.g. Lambrides et al., 2024; Inayoshi et al., 2024; Maiolino et al., 2025; Rusakov et al., 2025). As demonstrated in detail in Naidu et al. (2025), the combination of such a ‘black hole star’ (BH*) with a (very weak) host galaxy, to make up the rest-frame UV emission, reproduces the observed spectrum of A2744-QSO1 quite well. This scenario also perfectly explains our simultaneous detection of broad-line EW variability, which come from the BH*, and non-detection of rest-frame UV variability, which would originate from young stars forming in the vicinity of the BH*.
@@
Our results further show that robust investigations into the variability of LRDs, including A2744-QSO1, will require dedicated monitoring with numbers of epochs large enough to sample the light-curve frequently and maximize the probability to detect significant variation events. More importantly, we stress that such monitoring campaigns need to be consistent in depth, coverage, and observing strategy to minimize the systematics between single epochs. Thanks to the gravitational time delays, this source is ideally suited for further study of its variability and could make for a prime target for a large-scale reverberation mapping campaign in the future. With that, one could potentially measure the black-hole mass of an AGN at directly and, for the first time, calibrate the scaling relations used to measure black-hole masses beyond the near-by Universe (e.g. Liu et al., 2024). Thanks to SL time delays, multiply-imaged, cluster-lensed AGN like A2744-QSO1 may open a door to achieve this on a reasonable timescale out to high-redshifts and high black-hole masses (e.g. Golubchik et al. 2024).